OpenStudy (anonymous):

Starting with the equation, Ek = mv^2/2, find an equation for a) m and b) v.

OpenStudy (anonymous):

you mean just solve for m and v? should be m=(2Ek)/v^2 v=sqrt(2Ek/m)

OpenStudy (anonymous):

a) 2E/v^2 = m b) Square root(2E/m)

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

np